OpenStudy (anonymous):

need some help with this yucky problem :) from a ship offshore, the angle of elevation of a hill is 1.1 degrees. After the ship moves inland at 4.5 knots for 20 min, the angle of elevation is 1.4 degrees. How high is the hill? (1 knot = about 6080 feet per hour) i found out that the boat with the 1,1 degree angle is 9120 ft from its second position with the 1.4 degree angle, but now im stuck :(

OpenStudy (anonymous):



OpenStudy (anonymous):

there might be a snappier way to do this but i came up with this one. you know \[\tan(1.1)=\frac{h}{x+9120}\] and also \[\tan(1.4)=\frac{h}{x}\]

OpenStudy (anonymous):

we can solve this equation for x and then for h we have \[x\tan(1.4)=(x+9120)\tan(1.1)\] \[x\tan(1.4)=x\tan(1.1)+9120\tan(1.1)\] \[x\tan(1.4)-x\tan(1.1)=9120\tan(1.1)\] \[x(\tan(1.4)-\tan(1.1))=9120\tan(1.1)\] \[x=\frac{9120\tan(1.1)}{\tan(1.4)-\tan(1.1)}\]

OpenStudy (anonymous):

i get \[x=33428\] rounded. now \[\tan(1.4)=\frac{h}{33428}\] so \[h=33428\tan(1.4)=817\] rounded

OpenStudy (anonymous):

i did not check the number 9120, i took your word for it

OpenStudy (anonymous):

haha im think i got that part right, thanks! :D

OpenStudy (anonymous):

yw